### v2

# Quantum thermodynamic derivation of the energy resolution limit in magnetometry

Abstract
It was recently demonstrated that a large number of magnetic sensing technologies satisfy the energy resolution limit, which connects a quantity composed by the variance of the magnetic field estimate, the sensor volume and the measurement time, and having units of action, with $\hbar$. A first-principles derivation of the energy resolution limit is still elusive. We here present such a derivation based on quantum thermodynamic arguments. We show that the energy resolution limit is a result of quantum thermodynamic work associated with quantum measurement and Landauer erasure, the work being exchanged with the magnetic field. We apply these considerations to atomic magnetometers and SQUIDS. Regarding the former, we unravel a new spin correlation effect relevant to the magnetic noise produced by atomic vapors.
Magnetic fields convey useful information in diverse physical settings, therefore it is no surprise that quantum sensing [1] of magnetic fields is one of the pillars of the second quantum revolution [2]. Whether classical or quantum, magnetometers are characterized by several figures of merit like bandwidth [3, 4, 5, 6], dynamic range [7, 8, 9, 10], sensor size and scalability [11, 12, 13, 14], accuracy [15], or the ability to operate in harsh environments [16, 17].
Magnetic sensitivity stands out as a prominent sensor characteristic, since the resolution of any kind of measurement is limited by the intrinsic noise properties of the sensor. Advances in magnetic sensitivity brought about by superconducting sensors [18, 19] and optical pumping magnetometers [20, 21] have spurred numerous applications, like sensing magnetic fields produced by the human brain [22, 23, 24, 25, 26] or heart [27, 28], materials characterization [29, 30], even developing table-top probes of new physics [31, 32, 33, 34, 35]. Therefore, understanding the fundamental limitations to magnetic sensitivity is crucial for pushing any kind of magnetic sensor towards optimal performance, and thus unraveling new applications.
Using an analysis of a large body of published work, the authors in [36] demonstrated that tens of different magnetic sensing technologies appear to have a unifying property, namely they all seem to satisfy the so-called energy resolution limit. This limit states that $(\delta B)^{2}V\tau/2\mu_{0}\gtrapprox\hbar$, where $(\delta B)^{2}$ is the variance of the magnetic field estimate, $V$ the sensor volume, $\tau$ the measurement time, and $\mu_{0}$ the magnetic permeability of the vacuum. The left-hand side of this bound has units of action. The fact that the right-hand side roughly equals $\hbar$ is aesthetically pleasing when discussing a fundamental limit. Surprisingly, as the authors in [36] pointed out, a first-principles derivation of the energy resolution limit (ERL) is still elusive.
The ERL contains the expression $(\delta B)^{2}V/2\mu_{0}$, which is reminiscent of the magnetic energy within the sensor volume $V$. However, the expression $(\delta B)^{2}/2\mu_{0}$ is not the actual magnetic energy density, since it contains $(\delta B)^{2}$ instead of $B^{2}$. The authors in [36] also showed that various attempts to derive the ERL based on quantum speed limits or energy-time uncertainty relations do not work, since instead of $(\delta B)^{2}$, they involve the expression $\delta(B^{2})=2B\delta B$, further leading to the counterintuitive result that $\delta B$ is suppressed when increasing $B$.
We will here derive the ERL based on quantum thermodynamic arguments. In particular, we will connect the ERL to the quantum work performed during quantum measurement and/or Landauer erasure of information. To this end, we treat the magnetic field as an integral part of the quantum thermodynamic environment of the sensor. We will show that the quantum thermodynamic work accompanying the process of measurement is taken up by the magnetic field energy, and thus leads to magnetic field fluctuations.
The field of quantum thermodynamics [37, 38, 39, 40, 41, 42, 43] has unified quantum information and quantum measurements with thermodynamic processes. The understanding of the physical nature of information [44, 45], and the energy cost of information erasure [46] greatly inspired the development of quantum information science. Few works, however, have so far touched upon the connection of quantum thermodynamics with quantum metrology [47, 48, 49]. The current work falls in this direction.
The idea that the magnetic or electric field itself can act as a source/sink of quantum thermodynamic work is not new [47, 50]. We here push this idea further, towards understanding the ERL in magnetometry in a general way independent of the specific technology realization. The crux of the matter is the following. Let $u_{B}=B^{2}/2\mu_{0}$ be the magnetic energy density produced by the magnetic field to be sensed. If the field fluctuates by $\delta B$, where $\braket{\delta B}=0$, then the field energy within the volume $V$ of the sensor will change by $V\braket{u_{B+\delta B}-u_{B}}=(\delta B)^{2}V/2\mu_{0}$. If the cause of this fluctuation is the exchange of work $W$ between sensor and field, it will be $W=V\braket{u_{B+\delta B}-u_{B}}$, and the corresponding field fluctuation will be $\delta B=\sqrt{2\mu_{0}W/V}$. Finally, from the relation $W=(\delta B)^{2}V/2\mu_{0}$, and by using quantum speed limits to connect the exchanged work $W$ and the time during which the exchange takes place, we will arrive at the ERL.
To proceed formally and find the value of $W$ we will use the approach of [51], by which the authors establish the minimum energy cost of measurement and Landauer erasure. The authors consider a system ${\cal S}$, a memory ${\cal M}$, and a thermal bath ${\cal B}$. A measurement is performed on ${\cal S}$, meaning that ${\cal S}$ and ${\cal M}$ become entangled. Then follows a projective measurement on ${\cal M}$. Finally, to make the process cyclic, the information in ${\cal S+M}$ is erased. The authors show that the combined work cost (done by ${\cal M}$ on ${\cal B}$) of measurement and erasure is bounded below:
$W_{\rm meas}+W_{\rm eras}\geq k_{B}T{\cal I},$ (1)
where $T$ is the bath temperature, with which the memory is in thermal contact, and ${\cal I}$ the mutual information between ${\cal S}$ and ${\cal M}$. This information satisfies $0\leq{\cal I}\leq H$, where $H$ is the Shannon entropy of the possible measurement results. If this work is exchanged between sensor and magnetic field, it will be $(\delta B)^{2}V/2\mu_{0}=k_{B}T{\cal I}$. If the exchange takes place during time $\tau$, we can use the Margolus-Levitin quantum speed limit [61, 62], $k_{B}T{\cal I}\tau\geq{\pi\over 2}\hbar$, and finally we arrive at the ERL:
${{(\delta B)^{2}V\tau}\over{2\mu_{0}}}\geq{\pi\over 2}\hbar$ (2)
We will now specify the above for the case of optical pumping magnetometers, so the system ${\cal S}$ doing the sensing will be an atomic vapor. Hyperfine structure is not relevant in this discussion, so we consider a vapor of spin-1/2 atoms. We consider the standard magnetic sensing framework where the atomic spins are first spin-polarized by an optical pumping pulse, then precess under the action of the magnetic field to be sensed, and finally they are probed e.g. by a light beam. The computational basis states |$\ket{0}$ and |$\ket{1}$ are eigenstates of $\sigma_{z}$, and the magnetic field to be sensed is along the $z$-axis, $\mathbf{B}=B\mathbf{\hat{z}}$. The atoms are initially spin-polarized along the $x$-axis, so their initial state is |$\ket{\psi_{0}}=(\ket{0}+\ket{1})/\sqrt{2}$. The magnetic field induces a relative phase $\chi$ between |$\ket{0}$ and |$\ket{1}$, i.e. it will produce the state |$\ket{\psi_{\chi}}=(\ket{0}+e^{i\chi}\ket{1})/\sqrt{2}$, where $\chi\propto B$. A measurement in the eigenbasis of $\sigma_{x}$, e.g. through the interaction of the atoms with a probe laser, conveys information about $B$.
However, in our analysis, as measurement in the quantum thermodynamic context of [51] we will consider the dephasing produced by atomic collisions. In particular, in the SERF regime of interest here, spin destruction (SD) collisions [52] are the fundamental mechanism for spin decoherence. In a binary SD collision one atom is the system ${\cal S}$ and another atom is the memory ${\cal M}$. During the collision, the two atomic spins become entangled, and further SD collisions act as a projective measurement on the memory atom, along the same lines described in [53] for the case of spin-exchange collisions. Thus, SD collisions can be seen as performing an unobserved measurement of |$\ket{\psi_{\chi}}$ in the computational basis, pushing the state |$\ket{\psi_{\chi}}$ towards $\rho={1\over 2}\ket{0}\bra{0}+{1\over 2}\ket{1}\bra{1}$ |. The measurement time $\tau$ is the duration of this process, given by the spin-relaxation time attributed to SD collisions.
Information erasure (e.g. by an optical pumping pulse) renders the whole process cyclic. Nevertheless, information erasure does not incur an energy cost, since the optical pumping photons are scattered into the light field having practically zero temperature [54]. Hence in our case, the thermodynamic work entering (1) is solely due to the aforementioned measurement. Finally, the Shannon entropy of the decohered state $\rho$ is $H=\ln 2$. We still need to determine the mutual information ${\cal I}$. When a system atom collides with a bath atom, their spins become entangled. The entanglement is not maximal, because the atom’s spin phase change in SD collisions is small [55], unlike the case of spin-exchange collisions [56, 53]. Hence it might appear that ${\cal I}\ll H$. However, after many binary collisions between the system atom and bath atoms, the system state is fully decohered. We consider all those collisions as one process mapping |$\ket{\psi_{\chi}}$ into $\rho$ along the measurement time $\tau$, during which the total mutual information ${\cal I}=H$.
To apply the bound $k_{B}T{\cal I}\tau\geq{\pi\over 2}\hbar$, we crucially note that the temperature entering the expression $k_{B}T$ should be the spin temperature $T_{s}$, which can be widely different than the usual thermodynamic temperature. While the translational degrees of atoms are indeed governed by the latter, which is the room temperature or higher, the spin temperature is usually much lower. To find $T_{s}$ when the magnetic field is $B$ we relate $\mu B/k_{B}T_{s}$ with the spin-noise polarization of the vapor along the $z$ axis, which is on the order of $1/\sqrt{N}$, where $N$ is the atom number, and $\mu$ the atom’s magnetic moment. Thus $k_{B}T_{s}=\mu\sqrt{N}B$. This point is further substantiated by the fact that an atom in the coherent superposition state |$\ket{\psi_{\chi}}$ has zero energy, given that the atom’s Hamiltonian in the magnetic field $\mathbf{B}$ is proportional to $\sigma_{z}$. When the atom’s state is finally projected to either |$\ket{0}$ or |$\ket{1}$ due to SD collisions, the atom acquires energy ±$\pm\mu B$, which is provided by the field. Since the probabilities for projection to either |$\ket{0}$ or |$\ket{1}$ are equal, the total exchanged energy with the field is on average zero, apart from fluctuations of order $\sqrt{N}\mu B$, which is the same quantity derived previously as $k_{B}T_{s}$.
Setting $B\approx\delta B$, i.e. working at small magnetic fields close to the sought after noise, and using the bound $k_{B}T_{s}{\cal I}\tau\geq\pi\hbar/2$, we find $\delta B\geq(\pi/2\ln(2))(\hbar/\sqrt{N}\tau)$. The relaxation time due to SD collisions is given by $1/\tau=n\sigma_{\rm sd}\overline{v}$, where is $n=N/V$ the atom number density, $\sigma_{\rm sd}$ the SD cross section, and $\overline{v}$ the relative velocity of the colliding partners. Hence
$\delta B\geq{\pi\over{2\ln 2}}{{\hbar\sigma_{\rm sd}\overline{v}\sqrt{N}}\over{\mu V}}$ (3)
The atom’s magnetic moment is $\mu=\mu_{B}/q$, where $\mu_{B}$ is the Bohr magneton and $q=[(S(S+1)+I(I+1)]/S(S+1)$ is the nuclear slowing down factor with $S=1/2$ and $I$ the nuclear spin [52]. Using the SD cross sections listed in [52] and the respective relative velocities $\overline{v}$ we find for a number density $n=10^{14}~{}{\rm cm^{-3}}$ and volume $V=10~{}{\rm cm^{3}}$ that $\delta B=2\times 10^{-17}~{}{\rm T}$ for ${}^{41}{\rm K}$, $\delta B=10^{-16}~{}{\rm T}$ for ${}^{87}{\rm Rb}$, and $\delta B=10^{-14}~{}{\rm T}$ for ${}^{133}{\rm Cs}$. Using instead the parameters of [57], where the authors used a cesium vapor of volume $V=1~{}{\rm cm^{3}}$ and number density $2\times 10^{13}~{}{\rm cm^{-3}}$, we find again $\delta B\approx 10^{-14}~{}{\rm T}$. Since this noise is distributed within the bandwidth $1/\tau$, the corresponding spectral density is $10~{}{\rm pG/\sqrt{Hz}}$, whereas the authors measured a noise level of $400~{}{\rm pG/\sqrt{Hz}}$, and under the premise of a theoretical optimization project a noise level of $2~{}{\rm pG/\sqrt{Hz}}$. Finally, using (3) we can obtain the ERL which reads
${{(\delta B)^{2}V\tau}\over{2\mu_{0}}}\geq\hbar\Big{(}{{\pi^{2}}\over{8(\ln 2)^{2}}}{{\hbar\sigma_{\rm sd}\overline{v}}\over{\mu_{0}\mu^{2}}}\Big{)}$ (4)
The ERL is $4\hbar$ for ${}^{41}{\rm K}$, $25\hbar$ for ${}^{87}{\rm Rb}$, and $6054\hbar$ for ${}^{133}{\rm Cs}$.
We will now elaborate further on the bound (3), rewriting it as
$\delta B\geq{2\over 3}{\kappa\mu_{0}\mu\sqrt{N}\over V},$ (5)
in order to exhibit the magnetic field produced within the magnetized volume of a randomly spin-polarized vapor. Only now, this dipolar field is seen to be amplified by the factor $\kappa=({{3\pi}/{4\ln 2}}){{\hbar\sigma_{\rm sd}\overline{v}}/{\mu_{0}\mu^{2}}}$. On the one hand we can argue, like in [36], that this amplification is forced by the uncertainty relation. Given that the uncertainty $\Delta\sigma_{x}$ is negligible for a vapor optically pumped along the $x$-axis, the uncertainty $\Delta\sigma_{y}$ will be determined by the field $\delta B$ driving spin precession from the $x$ axis to the $y$-axis, i.e. $\Delta\sigma_{y}=\delta\phi\braket{\sigma_{x}}$ where $\delta\phi=\mu\delta B\tau/\hbar$. Since $\Delta\sigma_{y}$ and $\Delta\sigma_{z}$ have to satisfy the uncertainty relation $\Delta\sigma_{y}\Delta\sigma_{z}\geq|\braket{\sigma_{x}}|$, it follows by setting $\Delta\sigma_{z}=1/\sqrt{N}$ that $\delta B$ is given by (5). The expectation value $\braket{\sigma_{x}}$ drops out in the previous calculation, so the same argument holds for an unpolarized vapor.
There is yet another interpretation for the factor $\kappa$ in (5), interesting in its own right. We can consider the vacuum permeability $\mu_{0}$ is replaced by $\kappa\mu_{0}$, as if the spins have a tendency to align, and $\kappa$ represents the relative permeability constant. But indeed, binary spin-dependent collisions do have the tendency to correlate atoms, as was shown in [53] for the case of spin-exchange collisions. As mentioned before, spin-destruction collisions are described by a spin phase change $\phi$, which is $\phi\ll 1$. In [53] it was shown that the negativity of the two-atom spin state scales as $\sin\phi_{\rm se}^{2}$, where $\phi_{\rm se}\approx 1$ is the spin-exchange phase change. Similarly, here we can state that the entanglement between two colliding atoms will scale like $\sin\phi^{2}\approx\phi^{2}$. Hence any two colliding partners will share a small correlation. The upside is that this correlation can be shared by many more atoms. Imagine a ”bath” atom experiencing consecutive collisions with many ”system” atoms. These will all be slightly correlated, and will reside in a ”correlation volume” $V_{c}$. Consider two spins inside this volume, which are about to collide. Their interaction energy is $\epsilon=\kappa\mu_{0}\mu^{2}/\tilde{V}$, where $\tilde{V}$ is the volume defined by the two spins. This interaction reorients the spins in a time scale $\hbar/\epsilon$. For the aforementioned correlation to be maintained, this time should be equal or larger than $\tilde{V}/\sigma_{\rm sd}\overline{v}$, which is the time required for their collision to correlate them. From this requirement we obtain again $\kappa=\hbar\sigma_{\rm sd}\overline{v}/\mu_{0}\mu^{2}$. We can also estimate the correlation volume $V_{c}$, which will contain $N_{c}$ atoms. We can write $\kappa=\sqrt{N_{c}}\phi$, since the spin variance scales as $\phi^{2}$ and thus the uncertainty as $\phi$, while for $\phi\approx 1$ we would get an enhancement $\kappa=\sqrt{N_{c}}$, retrieving the case of fully polarized spins. The phase $\phi$ is related to the SD relaxation time and the collision rate $1/T$, i.e. it is $1/\tau=\phi^{2}/T$ as discussed in [56]. The collision rate is $1/T=\overline{v}/n^{-1/3}$. Putting everything together we find $N_{c}=(\hbar\overline{v}/\mu_{0}\mu^{2})^{2}\sigma_{\rm sd}n^{-2/3}$. For a ${}^{41}{\rm K}$ number density of $n=10^{14}~{}{\rm cm}^{-3}$ it is $N_{c}\approx 10^{9}$, and $V_{c}=N_{c}/n\approx 0.01~{}{\rm mm}^{3}$. For ${}^{133}{\rm Cs}$ these numbers would be $N_{c}\approx 10^{13}$, and $V_{c}\approx 0.1~{}{\rm cm}^{3}$.
The physical picture that emerges is that both a polarized vapor being relaxed, as well as an unpolarized vapor, will behave like a “squashy” spin medium exhibiting spin fluctuations, split in a number of “domains” not unlike ferromagnetic systems. Such domains exhibit intra-domain correlations, but not inter-domain correlation. The vapor as a whole continuously exchanges energy with its self-field, setting an unavoidable noise level for measuring externally applied magnetic fields. Towards an experimental verification of the noise $\delta B$, we propose the use of a miniaturized cesium cell of volume e.g. $1~{}{\rm mm^{3}}$, in order to boost the previous estimates of $\delta B$ by a factor of $10^{3}$. The cell should be surrounded by a superconducting flux transformer in order to alleviate Johnson noise that would dominate the signal induced by the changing flux produced by $\delta B$ in an ordinary coil.
The possibility that spin uncertainty in an atomic vapor creates a fluctuating field consistent with the ERL has been discussed in [36] and references therein, but largely dismissed, mainly on grounds of angular momentum and energy conservation. The physical picture painted here is in several subtle ways different from the discussion in [36]: (i) We claim that when starting from a non-equilibrium i.e. spin-polarized state, which during the time $\tau$ relaxes to the equilibrium state exhibiting spin-noise fluctuations, there will be an energy exchange between atomic spins and magnetic field. (ii) Based on the understanding obtained from our physical description, we claim that this exchange will continue in subsequent time intervals $\tau$ even if the atoms are left alone in their equilibrium state. That is, while the authors in [36] refer to quantum measurement induced spin uncertainty, we refer to spontaneous spin noise that exists in such vapors whether there is or isn’t an external measurement limited by quantum uncertainty. (iii) There is no issue with angular momentum conservation brought up in [36], since in SD collisions angular momentum is anyhow taken up by translational angular momentum of the colliding atoms. Similarly, there is no issue with energy conservation, since spin dynamics are driven by collisions, the translational energy of which is supplied by an external energy source. A tiny fraction of this energy is transformed into magnetic interactions through SD and other spin-dependent collisions. (iv) While the authors in [36] postulated such magnetic field fluctuations emanating from spin uncertainty, and showed they are consistent with the ERL, we started from the opposite direction. We postulated the thermodynamic work exchanged between sensor and field, and arrived at these spin fluctuations and their self-interaction, connecting the thermodynamic work to the specific measurement process going on during SD spin relaxation. (v) Additionally, the authors in [36] considered two different noise sources, one stemming from the uncertainty relation discussed above, and one stemming from the field produced by randomly polarized spins. Then they elaborated on the difficulties of adding those two terms in quadrature. In contrast, we unify these into a single noise term, which contains within it the issue of the uncertainty relation, also interpreted with the novel type of correlations we unravel as an enhanced permeability of the vapor. (vi) The quantum speed limits did not work in the considerations of [36], because they concerned the total energy $B^{2}V/2\mu_{0}$. In our case, they concern the energy exchanged between atoms and field, which is $(\delta B)^{2}V/2\mu_{0}$.
A pending question is whether the ERL is a hard fundamental limit. In fact, there are works [64, 65, 66] claiming measurements of the ERL below $\hbar$. We note that the magnetic field fluctuations in (3) were derived assuming the equilibrium state of the atomic vapor exhibits spin fluctuations of order $\sqrt{N}$, pertinent to the standard quantum limit of uncorrelated atoms. If this is not the case, i.e. if for example the equilibrium state of the spin system happens to exhibit correlations, then the spin fluctuations will be suppressed by some squeezing factor $\xi\leq 1$ [67]. Correspondingly, the ERL in (4) will be violated by the factor $\xi^{2}$. In this respect, the claims in [64, 65, 66] seem plausible. Given that the Heisenberg limit allows $\xi^{2}=1/N$, where $N$ is the atom number, one could imagine an ERL at the ultra-low level of $\hbar/N$. The previous conclusions tacitly assume that the relaxation time $\tau$ does not show a $\xi$-dependence. However, the physics of spin relaxation in correlated vapors is currently not understood, although some works have touched upon it [68, 69, 70]. Thus at this point, and notwithstanding the claims [64, 65, 66], we cannot make strong statements about the general validity of the ERL.
We will, however, provide yet another demonstration of the validity of the ERL for a different sensor technology, the SQUID. Going back to the relation $(\delta B)^{2}V/2\mu_{0}=k_{B}T{\cal I}$, the fluctuation $\delta B$ will be $\delta B=\sqrt{2\mu_{0}k_{B}T{\cal I}/V}$, but now we set $T\approx 4~{}{\rm{}^{\circ}K}$. What is the information ${\cal I}$ extracted during a measurement with the SQIUID? The SQUID sensor could be seen as distinguishing between two alternatives, a flux equal to $n\Phi_{0}$ and a flux equal to $(n+1)\Phi_{0}$, where $\Phi_{0}$ is the flux quantum. These two choices would indeed correspond to 1 bit of information. However, in the flux-locked-loop operating mode [18, 19], the actual flux noise (the excursions of the flux around the operating point $(n+1/2)\Phi_{0}$) is on the order of $p\Phi_{0}$, so the gained information is thus ${\cal I}\approx-p\ln p$, where $p\ll 1$. For example, using the parameters of [63], namely a coil area $A=10^{-3}~{}{\rm m^{2}}$, thus an effective volume $V=A^{3/2}$, and flux noise $0.6\times 10^{-6}\Phi_{0}$, we obtain $\delta B\approx 6~{}{\rm fT}$, consistent with the average noise observed in [63] within the bandwidth. Moreover, the measurement time in [63] was $\tau=10^{-5}~{}{\rm s}$. It follows that $(\delta B)^{2}V\tau/2\mu_{0}=k_{B}T{\cal I}\tau=45\hbar$, whereas the authors estimate $35\hbar$.
Concluding, we have presented a first-principles derivation for the energy resolution limit in magnetic sensing, specified for atomic spin magnetometers and SQUIDs. The derivation was based on quantum thermodynamic arguments connecting magnetic field fluctuations to quantum thermodynamic work associated with quantum measurement. We obtain realistic values for magnetic noise by applying the same main principle for starkly different physical parameter values relevant to two starkly different sensor technologies. For atomic magnetometers in particular, we demonstrate the central role of spin noise, and unravel interesting effects related to collision-induced correlations acting to enhance the magnetic permeability of small spin-correlated “domains”.